OpenStudy (anonymous):

cos θ - cos θ sin^2θ

OpenStudy (anonymous):

\[\cos θ - \cos θ \sin^2θ\]\[\cos θ(1 - \sin^2θ)\]\[\cos θ(\cos^2θ)\]\[\cos^3θ\]

OpenStudy (anonymous):

how do you do the second part? cosθ(1−sin2θ)

OpenStudy (anonymous):

I used the rule, \[\sin^2θ+\cos^2θ=1\]from this rule, if you subtract sin²θ from both sides, you get \[1-\sin^2θ=\cos^2θ\]

OpenStudy (anonymous):

ohhhh right got it thank you